solve indefinite integral from -1 to 1 (1-x)dx?
1
Expert's answer
2012-03-02T07:15:44-0500
First of all, as long as you asked to "solve integral from -1 to 1 " it's definite integral.
Integral(1-x)dx|-11={x-x2/2}|-11 =(1-1/2)-(-1-1/2)=1-1/2+1+1/2=2